David is the age of 58 and Lisa is the age of 18. How long ago was David 5 times the age of Lisa.
Logically, I know the answer is 8 years ago since David would have been 50 and Lisa would have been 10 and 5*10 =50 which five times her age.
However, I need to come with this answer in an algebraic method and I've been wracking my brain over it.

Answer by ankor@dixie-net.com(22740)

David is the age of 58 and Lisa is the age of 18.
How long ago was David 5 times the age of Lisa.
:
let y = the no. years ago that D was 5 times L's age
:
58 - y = 5(18-y)
58 - y = 90 - 5y
-y + 5y = 90 - 58
4y = 32